# Optimizing wheel loader performance: an end-to-end approach

### v1

@@ 4.4.1 Greedy strategy
We refer to the choice of $d=1$ as an entirely greedy strategy. The dig location is always selected to optimize the one-step evaluation function $Q^{d=1}_{n}=Q_{\text{greedy}}\equiv\bm{w}^{\mathsf{T}}\bm{\mathcal{P}_{n}}$ with the loading action optimized for the local pile state at the selected dig location. In this case, the pile state $\bm{H}_{n}$ needs only to be expanded once and not into multiple pile state candidates.
@@ 4.5 Computational time
The computational time is what limits us in searching for an optimal solution. We register the computational time for analyzing a single loading cycle. It includes predicting the new pile state $\bm{H}^{\prime}$ and total performance $\bm{\mathcal{P}}$, which involves updating $\bm{a}^{\text{load}}$ accorging to Eq. (9) (dominant part) and computing and integrating the V-turns given a selected dig location $\bm{x}^{\text{dig}}$ as described in Sec. 4.2. Table 1 shows the average computational time from profiling the loading predictions made on a desktop computer with an Intel(R) Core(TM) i7-8700K, 3.70 GHz, 32 GB RAM on a Windows 64-bit system and NVIDIA GeForce RTX 2070 SUPER.
@@ 5 Results
The initial pile surface $\bm{H}_{1}$ features a trapezoidal prism, 1.8 m tall with a $30^{\circ}$ slope at the front, with added Perlin noise [16]. A representative initial pile surface is illustrated in Fig. 6.
@@
The location and orientation of the receiver are fixed at $\bm{x}^{\text{dump}}=[x,y,\theta]=[-12.0,-3.0,-30.0^{\text{o}}]$. The dig locations are constrained within $-5.0\leq x\leq 8.0$ and $0.0\leq y\leq 6.0$. The listup function generates candidate dig locations $\tilde{\bm{x}}^{\text{dig}}$ with a spacing of 1 m.
@@ 5.1 Greedy strategy
The greedy strategy with no horizon, $d=1$, is tested using the three evaluation functions described in Sec. 4.4.1: $Q_{\text{greedy}}=\bm{w}^{\mathsf{T}}\bm{\mathcal{P}}_{n}$, $Q_{\text{loading}}=\bm{w}^{\mathsf{T}}\bm{\mathcal{P}}_{n}^{\text{load}}$, and $Q_{\text{nominal}}=\bm{w}^{\mathsf{T}}\left(\bm{\mathcal{P}}_{n}^{\text{V1}}+\bm{\mathcal{P}_{n}}^{\text{V2}}\right)$. The pile states after five loading cycles are shown in Fig. 7. The greedy strategy results in loadings that aim for areas in the search region with more soil easily accessible from the dump truck. The maximum loading strategy aims to the right where there is more soil but further from the dump truck. The nominal strategy loads to the left, in proximity to the dump truck, as can be expected when focusing entirely on minimizing the transportation cost.
@@ 5.2 Long-horizon planning with look-ahead tree search
The second test uses the tree search algorithm with finite prediction horizon $d$ and evaluation function (4.4). Ten initial heightfields were first prepared with different Perlin noise applied to the 1.8 m high trapezoidal prism. For each of these ten piles, $N=15$ sequential loading cycles were tested with the prediction horizon ranging from $d=1$ to $d=N$. The resulting loading outcome and performance evaluation function are shown in Fig. 9. The results have been averaged over the ten repetitions starting with the different initial piles. The performance converges around $d=4$. Beyond this point, the total performance no longer improves despite the increased number of predictions executed in the search.
@@ 6 Discussion
With the look-ahead search method, it is possible to find action sequences that maximize the total performance over long horizons by increasing the pile state quality for future loadings. This method adapts to the truck’s location and can do so even if it would vary provided that the location is known in advance. In our experiments, the optimal action sequence can be (before any serious code optimization) computed in 43 s with a look-ahead search with depth $d=4$, which is close to the average loading cycle time. Future work should consider other soil properties, such as cohesive or inhomogeneous soil. The resulting pile state is likely to depend more strongly on the selected loading actions than it does for gravel, for which the pile always tends to settle in with a slope close to the natural angle of repose. Capturing this with a world model would be more demanding and might benefit from a longer horizon than $d=4$. That would motivate substantial improvements in the computational efficiency of the look-ahead search.
@@ Acknowledgement
Appendix A Example Appendix Section
Appendix text.

### v2

@@ 4.4.1 Greedy strategy
We refer to the choice of $d=1$ as the entirely greedy strategy. The dig location is always selected to optimize the one-step evaluation function $Q^{d=1}_{n}=Q_{\text{greedy}}\equiv\bm{w}^{\mathsf{T}}\bm{\mathcal{P}_{n}}$ with the loading action optimized for the local pile state at the dig location. In this case, the pile state $\bm{H}_{n}$ needs only to be expanded once and not into multiple pile state candidates.
4.4.2 Maximum loading strategy
@@ 4.5 Computational time
The computational time is what limits us from computing the optimal solution by exhaustive search. We register the computational time for analyzing a single loading cycle. It includes predicting the new pile state $\bm{H}^{\prime}$ and total performance $\bm{\mathcal{P}}$, which involves updating $\bm{a}^{\text{load}}$ accorging to Eq. (9) (dominant part) and computing and integrating the V-turns given a selected dig location $\bm{x}^{\text{dig}}$ as described in Sec. 4.2. Table 1 shows the average computational time from profiling the loading predictions made on a desktop computer with an Intel(R) Core(TM) i7-8700K, 3.70 GHz, 32 GB RAM on a Windows 64-bit system and NVIDIA GeForce RTX 2070 SUPER.
@@ 5 Results
The initial pile surface $\bm{H}_{1}$ features a trapezoidal prism, 1.8 m tall with a $30^{\circ}$ slope at the front, with Perlin noise [25] added to the surface. A representative initial pile surface is illustrated in Fig. 6.
@@
The location and orientation of the receiver are fixed at $\bm{x}^{\text{dump}}=[x,y,\theta]=[-12.0\text{ m},-3.0\text{ m},-30.0^{\text{o}}]$. The dig locations are constrained within 5.0 $-5.0\text{ m}\leq x\leq 8.0\text{ m}$ and 0.0 $0.0\text{ m}\leq y\leq 6.0\text{ m}$. The listup function generates candidate dig locations $\tilde{\bm{x}}^{\text{dig}}$ with a spacing of 1 m.
@@ 5.1 Greedy strategy
The greedy strategy with no horizon, $d=1$, is tested using the three evaluation functions described in Sec. 4.4.1-4.4.3: $Q_{\text{greedy}}=\bm{w}^{\mathsf{T}}\bm{\mathcal{P}}_{n}$, $Q_{\text{loading}}=\bm{w}^{\mathsf{T}}\bm{\mathcal{P}}_{n}^{\text{load}}$, and $Q_{\text{nominal}}=\bm{w}^{\mathsf{T}}\left(\bm{\mathcal{P}}_{n}^{\text{V1}}+\bm{\mathcal{P}_{n}}^{\text{V2}}\right)$. The pile states after five loading cycles are shown in Fig. 7. The greedy strategy results in loadings that aim for areas in the search region with more soil easily accessible from the dump truck. The maximum loading strategy aims to the right where there is more soil but further from the dump truck. The nominal strategy loads to the left, in proximity to the dump truck, as can be expected when focusing entirely on minimizing the transportation cost.
@@ 5.2 Long-horizon planning with look-ahead tree search
The second test uses the tree search algorithm with finite prediction horizon $d$ and evaluation function (4.4). Ten initial heightfields were prepared with different Perlin noise applied to the 1.8 m high trapezoidal prism. For each of these ten piles, $N=15$ sequential loading cycles were evaluated with the prediction horizon ranging from $d=1$ to $d=N$. The resulting loading outcome and performance evaluation function are shown in Fig. 9. The results have been averaged over the ten repetitions starting with the different initial piles. The performance converges around $d=4$. Beyond this point, increasing the search horizon further does not enhance the overall performance.
@@ 6 Discussion
With the look-ahead search method, it is possible to find action sequences that maximize the total performance over long horizons by increasing the pile state quality for future loadings. This method adapts to the truck’s location and can do so even if it would vary provided that the location is known in advance. In our experiments, the optimal action sequence can be (before any serious code optimization) computed in 43 s with a look-ahead search with depth $d=4$, which is close to the average loading cycle time. Future work should consider other soil properties, such as cohesive or inhomogeneous soil. The resulting pile state is likely to depend more strongly on the selected loading actions than it does for gravel, for which the pile always tends to settle with a local slope near the natural angle of repose. Capturing this with a world model would be more demanding and might benefit from a longer horizon than $d=4$. That would motivate substantial improvements in the computational efficiency of the look-ahead search.